5 10 15 20 25 30 35 40

illustrate the variance and standard deviation


1
Expert's answer
2022-04-26T10:32:48-0400

"\mu =(5+10+15+20+25+30+35+40)/8=22.5"

"D=\frac{\sum (x-\mu)^2}{n}=\frac{306.25+156.25+56.25+6.25+6.25+56.25+156.25+306.25}{8}=131.25"

"\sigma=\sqrt D=\sqrt{131.25}=11.46"